For a Binomial random variable, the probability of exactly zero successes out of two trials equals 0.0289. What is the associated probability of "success," p?
Answer #1

P[X=0] = C(2,0)*(P)0 * (1-P)2 = 0.0289

Therefore,

(1-P)2 = 0.0289

1-2P+P2 = 0.0289

P =1.17 or P=0.83

P= 1.17 is not possible

Therefore, p=0.83

***please ask if you have any doubts.Happy to help you.Thank you.Please Like.